OpenStudy (anonymous):

square root of 50 + square root of 6

OpenStudy (anonymous):

\(\Large \sqrt{50}+\sqrt{6}=\sqrt{25\times 2}+\sqrt{6}=5\sqrt{2}+\sqrt{6}\)

OpenStudy (anonymous):

Thank you @JungHyunRan

OpenStudy (anonymous):

^^ yes welcome